OpenStudy (anonymous):

If x / y is an integer, which of the following statements need NOT always be true? Select ALL such statements. A. both x and y are integers B. x is an integer C. either x or y is negative D. y / x is an integer E. x = ny where n is an integer

OpenStudy (bibby):

Is 1.2/1.2 an integer?

OpenStudy (anonymous):

I think so because that equals 1

OpenStudy (bibby):

3.4/3.4 = 1 = an integer 7/3.5 = 2 = an integer Unsure about C. D is not necessarily true. 2/1 is an integer but 1/2 is not. Can I just pick n=1 for every case of E?

OpenStudy (bibby):

gotta go bye ♥